OpenStudy (anonymous):

Find the values of c that satisfy the equation f(b)-f(a)/b-a= f prime(c) in the conclusion of the mean value theorem for the given function and interval. f(x)=9x+9/x [1/9, 9]

OpenStudy (anonymous):

Due to having to type, my arithmetic may be off, but the steps should be clear. First solve for f(a) and f(b) f(a) = 9a + 9/a = 9(1/9 + 1)/(1/9) = 90 f(b) = 9b + 9/b = 9(9 + 1)/(9) = 10 Then solve for f'(c) = f(b) - f(a)/(b - a) = 10 - 90/(9 - (1/9) = -90 Then differentiate f(x): f'(x) = 9 - 9/x^2 Substitute c: f'(c) = 9 - 9/c^2 = -90 Solve the equation for c! I haven't checked for errors, but the steps should be clear.

OpenStudy (anonymous):

I am not getting 90 or 10...

OpenStudy (anonymous):

My window closed while answering. First solve f(b) then solve f(a) then solve b - a: Since f(x) = 9x + 9/x, f(a) = 9a + 9/a = 9(1/9) = 9/(1/9) = 1 + 9 = 10 Do the same for f(b). b - a would be 9 - 1/9 Then f'(c) would equal f(b) - f(a)/(b - a) Then find f'(x). f(x) = 9x + 9/x leads to f'(x) = 9 - 9/x^2 (check this) substitute c: f'(c) = 9 - 9/c^2, which equals f(b) - f(b)/(b - a) That should give you an equation for which you can then solve for c.

OpenStudy (anonymous):

9/(1/9)=81

OpenStudy (anonymous):

\[f(1/9)= 9(1/9)+\frac{ 9 }{ 1/9 }\]